McCurdy Co.'s Class Q bonds have a 12-year maturity, $1,000 par value, and a 12% coupon paid semiannually (6% each 6 months), and those bonds sell at their par value. McCurdy's Class P bonds have the same risk, maturity, and par value, but the p bonds pay a 12% annualcoupon. Neither bond is callable. At what price should the annual payment bond sell?
Answer #1

As the bonds sell for par value
yield to maturity with semi annual compounding=coupon rate=12%

yield to maturity with annual compounding=(1+12%/2)^2-1=12.3600%

Price of the bond=1000*12%/12.36%*(1-1/(1+12.36%)^12)+1000/(1+12.36%)^12=978.07